# P3H-25-058, TTP25-027

### v1

@@ Abstract
It is common practice to explain deviations between data and Standard-Model predictions by postulating new particles at the TeV scale ad-hoc. This approach becomes much more convincing, if one successfully embeds the postulated particles into a UV completion which addresses other conceptual or phenomenological shortcomings of the SM. We present a study of an SO(10) grand unified theory which contains scalar leptoquark fields employed to explain the “flavour anomalies” in $b\rightarrow s$ and $b\rightarrow c$ decays. We find that the additional degrees of freedom improve the renormalization group evolution of the SM parameters and may explain some of the observed fermion masses. In particular, the light leptoquarks modify the renormalization-group evolution of the Yukawa couplings such that successful bottom-tau unification becomes possible in a minimal SO(10) GUT with only a $126$-plet coupling to fermions.
P3H-25-058, TTP25-027
TeV-scale scalar leptoquarks motivated by B anomalies
improve Yukawa unification in SO(10) GUT
Xiyuan Gao  111xiyuan.gao@kit.edu, and Ulrich Nierste $^{\dagger}$222ulrich.nierste@kit.edu
Institute for Theoretical Particle Physics, Karlsruhe Institute of Technology (KIT),
Wolfgang-Gaede-Str. 1, D-76131 Karlsruhe, Germany
@@ 1 Introduction
However, it would be premature to claim that most minimal GUTs are unrealistic: The loophole is the assumption of a particle desert between the electroweak and GUT scales, usually deduced from naturalness criteria Georgi:1979md . The impact of the deviation of the desert picture on Yukawa unification is best studied in supersymmetric GUTs, in which the infrared (IR) theory is the Minimal Supersymmetric Standard Model (MSSM). The threshold correction to the matching relation between SM and MSSM Yukawa couplings can be enhanced when $\tan\beta$, the ratio of the vacuum expectation values (VEVs) of the two MSSM Higgs doublets, is large Hall:1993gn ; Carena:1993ag ; Carena:1999py ; Diaz-Cruz:2000nvf ; Girrbach:2011an ; Deppisch:2018flu . This fact changes the predictions of minimal GUT, and together with successful gauge coupling unification, has been viewed as indirect support for low-scale supersymmetry. However, the SUSY GUT is far from minimal. It requires many more physical particles than the non-SUSY one, and the huge number of physical degrees of freedom brings a new puzzle that perturbative expansion could fail Milagre:2024wcg ; Bajc:2016efj . Recently, the authors of Ref. Patel:2023gwt ; Shukla:2024bwf proposed a new idea, that the wrong $b-\tau$ mass relation in minimal SU(5) can be resolved by introducing a large mass hierarchy among the particles within the same scalar multiplet. This hierarchy requires that some particles lie far below the $M_{\text{GUT}}$, and suggests the desert picture together with the naturalness criterion should be reconsidered. The wrong fermion mass pattern does not necessarily falsify the minimal GUT; rather, it implies that SM alone cannot serve as a viable IR theory. Some scalar particles from the GUT sector need to be included in the light spectrum. In some cases, such particles also lead to successful gauge coupling unification; see, for instance, Ref. Dorsner:2005fq ; Bajc:2006ia ; Preda:2022izo ; Preda:2024vas .
@@
The state-of-the art is to postulate the required light LQs ad-hoc, which remains unsatisfactory until these particles are embedded into a meaningful theory addressing fundamental puzzles of the SM. Thus it is a natural idea to analyze whether the (multi-)TeV scale leptoquarks are beneficial to GUTs, as we do in this paper. Although LQs can arise naturally in many partial unification frameworks, such as Pati-Salam (PS) theories Pati:1974yy , their masses and interactions with fermions remain puzzling. If the LQs are vector bosons, their masses originate from spontaneous gauge symmetry breaking and are their mass is protected by gauge symmetry. However, if PS unification is realized at the multi-TeV scale, the LQ-fermion coupling structure needs to respect an approximate $U(2)^{n}$ flavor symmetry to evade the bounds from processes involving light flavors, such as $K_{L}\rightarrow\mu e$ Barbieri:2011ci ; Isidori:2012ts ; Barbieri:2012uh ; Blankenburg:2012nx ; Faroughy:2020ina ; Antusch:2023shi . Additional vector-like Fermions and/or extended gauge groups are typically needed to achieve TeV-scale partial unification while preserving $U(2)^{n}$ at low energies DiLuzio:2017vat ; Bordone:2017bld ; Greljo:2018tuh ; Calibbi:2017qbu ; Davighi:2022bqf . On the other hand, scalar LQs can naturally preserve the chiral $U(2)^{n}$ symmetry, but their masses are are unprotected and suffer from the same fine-tuning problem as the SM-Higgs boson PhysRevD.27.1601 . This conceptual puzzle becomes an explicit problem in a (partial) unification framework whose scale is much higher than a TeV. In our view, the LQ explanation for $B$ anomalies could become more convincing if LQs originate from a most minimal GUT. Their light masses and specific coupling structures could emerge as a consistency requirement for successful Yukawa unification, so that the benefits of successful unification and explanation of flavor anomalies outweigh the nuisance of additional unprotected scalar masses.
@@
The paper is structured as follows. In Section 2, we revisit the minimal ways to construct the SO(10) Yukawa sector. We explain why $126_{H}$ is preferred and discuss its shortcomings. Next, we discuss the LQ spectrum in $126_{H}$, specify the the LQ-fermion interactions, and summarize the effective operators relevant for $b\rightarrow c,s$ transitions. In the following subsection, we present an overview on how these LQs can address the $B$ anomalies and examine what additional constrains are imposed by SO(10). In Section 3, we analyze how TeV-scale LQs can improve the RG evolution for the Yukawa couplings between light scalars and third-generation fermions. We show how the $b-\tau$ mass relationship is improved and how the LQ-fermion couplings exhibit fixed point behaviors. Since addressing $B$ anomalies requires large flavor mixing angles that are unphysical in SM, we then propose a possible solution to this newly arising problem. In Section 4, we summarize our main findings and outline the further work needed to strengthen our idea. To demonstrate the model-independence of our results, we show in Appendix A that $b-\tau$ unification can also be achieved with different light LQ spectra.
@@ 2.1 The landscape
Firstly, $126_{H}$ is tailor-made for tiny neutrino masses. $126_{H}$ contains a vacuum singlet under $G_{\text{SM}}$ and can give the right-handed neutrino $\nu_{R}$ a high-scale mass term. Upon integrating out $\nu_{R}$, the dim-5 Weinberg operator Weinberg:1979sa can generate a tiny mass for the active neutrino living in $\ell_{L}$ Minkowski:1977sc ; Mohapatra:1979ia . In addition, $126_{H}$ contains two Higgs doublets with opposite $SU(2)_{R}$ isospin. As a result, one of the Higgses only couples to $\overline{Q_{L}}u_{R},\overline{L_{L}}\nu_{R}$ and the other one only couples to $\overline{Q_{L}}b_{R},\overline{Q_{L}}e_{R}$. A hierarchical ratio $m_{t}/m_{b}$ is now possible, as long as the ratio $\tan\beta$ of the two Higgs doublets VEVs is large. Furthermore, $Y_{126}$ is symmetric and can be assigned third-generation specific, which gives the desired hierarchy structure among the charged fermion masses of the three generations. Indeed, Eq (4) cannot account for quark mixing and is unlikely to yield the precise quantities for the first- and second-generation fermion masses. However, we do not think these shortcomings require extending the Yukawa sector immediately. If one neglects Yukawa couplings much smaller than unity, the light-flavor structure contains merely vanishing or unphysical observables. Large neutrino mixing is not problematic either, because the Yukawa couplings enter the left-handed neutrino mass matrix $M_{\nu_{L}}$ non-linearly. The relation $M_{\nu_{L}}\propto Y_{126}$ could break down even with higher-order corrections. Therefore, the $\mathcal{O}(1)$ predictions of $126_{H}$ are approximately correct. In the following, we refer to the theory whose Yukawa sector can be well-approximated by Eq (4) as ‘minimal SO(10) GUT’.
@@ 2.2 Leptoquarks in $126_{H}$
Here,c denotes the standard charge conjugation operator, $Q_{R}=(u_{R},d_{R}),~L_{R}=(\ell_{R},\nu_{R})$ are $SU(2)_{R}$ doublets, and flavor indices are implicit. Eq (8) shows $R$-type LQs couple to quarks and leptons with opposite chiralities, while the $S$-type LQs — $S_{3},S_{1}$ for left-handed fields and $\widehat{S}_{3},\widehat{S}_{1}$ for right-handed fields— are chirality-specific.
@@
$b\rightarrow c\tau\nu$ $(\overline{c}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma^{\mu}\nu_{L})$
@@
$(\overline{c}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma^{\mu}\nu_{L})$
@@
$b\rightarrow s\tau\tau$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma^{\mu}\tau_{L})$ - $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\tau}_{R}\gamma^{\mu}\tau_{R})$
$b\rightarrow s\nu\nu$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ -
@@
$b\rightarrow s\tau\tau$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{R}\gamma^{\mu}\tau_{R})$ $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\tau}_{L}\gamma^{\mu}\tau_{L})$
$b\rightarrow s\nu\nu$ - $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$
@@
In the energy regime far below 1 TeV, LQs in Eq (8) can be safely integrated out, and then the scalar-type effective operators with the form $(\overline{q}\ell)(\overline{\ell^{\prime}}q^{\prime})$ are generated at tree-level. We apply a Fierz transformation to convert these operators into the standard basis, as summarized in Table 1. These operators follow the general discussion shown in Ref. Dorsner:2016wpm and can also be directly inferred from the chirality structure of the LQ-fermion couplings. It is worth noting that the left-right symmetry is not manifest in the table, because $\nu_{R}$ and the top quark are not included. Moreover, the scalar- and tensor-type operators require sizable $R_{2}-(R_{2}^{\prime})^{c}$ and $S_{1}-(S_{1}^{\prime})^{c}$ (or $S_{1}-(S_{1}^{\prime\prime})^{c}$) mixing. Without such mixings, a single LQ in the generic basis cannot couple simultaneously to both $Q_{L}$ and $Q_{R}$, and thus leads to merely the vector-type operators.
@@ 2.3 Addressing the $B$ anomalies
TeV scale LQs are related to the long-standing anomalies observed in semi-leptonic B decays. The $S_{3}$ LQ contributes to all the $b\rightarrow c,s$ transitions with SM-like operators. If $S_{3}$ is the unique LQ responsible for the $R(D^{(*)})$ anomaly, $C_{L}^{\nu}$, the coefficient of $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ is fixed by $SU(2)_{L}$ invariance. The predicted value for $C_{L}^{\nu}$ is too large to be compatible with the current bounds on $b\rightarrow s\overline{\nu}\nu$ transitions. A simple way out is to combine $S_{3}$ with $S_{1}$, whose contributions to $C_{L}^{\nu}$ can partly cancel Crivellin:2017zlb ; Crivellin:2019dwb the one of the former LQ. Assuming a cancellation of about $60\%$ in $C_{L}^{\nu}$, and including the operators induced by $S_{1}$, $R(D^{(*)})$ can be consistently explained at $1\sigma$ level Crivellin:2025qsq . The constraints on $C_{L}^{\nu}$ can be relaxed since the recent Belle-II data with the inclusive tagging method indicates $\text{Br}(B\rightarrow K\overline{\nu}\nu)$ exceeds the SM prediction by a factor of 5.4. Yet $\text{Br}(B\rightarrow K^{*}\overline{\nu}\nu)$ still imposes a tight bound and necessitates the right-handed operator $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ generated by $\widetilde{R}_{2}$. If its coefficient $C_{R}^{\nu}$ takes a proper value, it can suppress the $B\rightarrow K^{*}$ amplitude while allowing a sizable $B\rightarrow K$ rate He:2023bnk . $SU(2)_{L}$ invariance also implies $b\rightarrow s\overline{\tau}\tau$ transitions. The current limit on $\text{Br}(B\rightarrow K\overline{\tau}\tau)$ is weak due the experimental difficulty in identifying $\tau$ leptons. Interestingly, if closing the $\tau$ loop and attaching it to an off-shell photon, the penguin diagram can induce $b\rightarrow s\ell\ell$ ($\ell=e,\mu$) via lepton flavor universal operators Crivellin:2018yvo . The coefficient $C_{9}^{U}\sim-1$ can account for the $B\rightarrow K\ell\ell$ anomalies without violating the bound from $R(K^{(*)})$ Alguero:2023jeh . Although $B_{s}-\bar{B}_{s}$ mixing constraints disfavor the best-fit value for $C_{9}^{U}$, moderate cancellation by additional operators containing right-handed quarks can relax this tension Crivellin:2015era ; Capdevila:2023yhq . Such operators can be generated by $\widetilde{S}_{1}$ or $\widetilde{R}_{2}$.
The $R_{2}$ LQ can also explain the $R(D^{(*)})$ anomaly with the $(\overline{c}_{R}\sigma^{\nu\nu}b_{L})(\overline{\tau}_{R}\sigma_{\nu\nu}\nu_{L})$ and $(\overline{c_{R}}b_{L})(\overline{\tau}_{R}\nu_{L})$ operators, and it does not lead to the $b\rightarrow s\overline{\nu}\nu$ transition. This solution requires an $\mathcal{O}(1)$ $R_{2}-(R_{2}^{\prime})^{c}$ mixing angle, which brings an inhomogeneous term to the RG equation for the $\tau$ lepton mass Fedele:2023rxb :
@@
The $S_{1}$ LQ is constrained to be as heavy as the GUT scale, because it always couples to a pair of quarks in the minimal GUT framework Patel:2022wya and induces proton decay. On the other hand, the diquark couplings of $S_{3},\overline{S}_{1},\widetilde{S}_{1}$ are absent at tree level. This absence is not accidental but a consequence of the $U(1)_{\text{PQ}}$ symmetry contained in Eq (4).

### v2

# Wolfgang-Gaede-Str. 1, D-76131 Karlsruhe, Germany

@@ Abstract
It is common practice to explain deviations between data and Standard-Model (SM) predictions by postulating new particles at the TeV scale ad-hoc. This approach becomes much more convincing, if one successfully embeds the postulated particles into a UV completion which addresses other conceptual or phenomenological shortcomings of the SM. We present a study of an SO(10) grand unified theory which contains scalar leptoquark fields employed to explain the “flavour anomalies” in $b\rightarrow s$ and $b\rightarrow c$ decays. We find that the additional degrees of freedom improve the renormalization-group (RG) evolution of the SM parameters. In particular, the light leptoquarks modify the RG evolution of the Yukawa couplings such that successful bottom-tau unification becomes possible in a minimal SO(10) GUT with only a $126$-plet coupling to fermions. If we amend the Yukawa interaction of the minimal one-generation model with a second fermion multiplet and small flavor-violating terms, we find the flavour violation in the leptoquark couplings growing with the RG evolution while it stays small in the Yukawa interaction of the SM Higgs boson. By employing mass splittings among the members of the 126 -plet one can increase the effect and obtain large flavor violation in leptoquark couplings from tiny perturbations at the GUT scale, because the flavour-conserving limit is an unstable initial condition for the RG equations.
preprint: P3H-25-058, TTP25-027
@@ 1 Introduction
However, it would be premature to claim that most minimal GUTs are unrealistic: The loophole is the assumption of a particle desert between the electroweak and GUT scales, usually deduced from naturalness criteria Georgi:1979md . The impact of the deviation of the desert picture on Yukawa unification is best studied in supersymmetric GUTs, in which the infrared (IR) theory is the Minimal Supersymmetric Standard Model (MSSM). The threshold correction to the matching relation between SM and MSSM Yukawa couplings can be enhanced when $\tan\beta$, the ratio of the vacuum expectation values (VEVs) of the two MSSM Higgs doublets, is large Hall:1993gn ; Carena:1993ag ; Carena:1999py ; Diaz-Cruz:2000nvf ; Girrbach:2011an ; Deppisch:2018flu . This fact changes the predictions of minimal GUT, and together with successful gauge coupling unification, has been viewed as indirect support for low-scale supersymmetry. However, the SUSY GUT is far from minimal. It requires many more physical particles than the non-SUSY one, and the huge number of physical degrees of freedom brings a new puzzle that perturbative expansion could fail Milagre:2024wcg ; Bajc:2016efj . Recently, the authors of Ref. Patel:2023gwt ; Shukla:2024bwf proposed a new idea, that the wrong $b-\tau$ mass relation in minimal SU(5) can be resolved by introducing a large mass hierarchy among the particles within the same scalar multiplet. This hierarchy requires that some particles lie far below the $M_{\text{GUT}}$, and suggests the desert picture together with the naturalness criterion should be reconsidered. The wrong fermion mass pattern does not necessarily falsify the minimal GUT; rather, it implies that SM alone cannot serve as a viable IR theory. Some scalar particles from the GUT sector need to be included in the light spectrum. In some cases, such particles also lead to successful gauge coupling unification; see, for instance, Ref. Dorsner:2005fq ; Bajc:2006ia ; Preda:2022izo ; Goto:2023qch ; Preda:2024vas .
@@
The state-of-the art is to postulate the required light LQs ad-hoc, which remains unsatisfactory until these particles are embedded into a meaningful theory addressing fundamental puzzles of the SM. Thus it is a natural idea to analyze whether the (multi-)TeV scale leptoquarks are beneficial to GUTs, as we do in this paper. Although LQs can arise naturally in many partial unification frameworks, such as Pati-Salam (PS) theories Pati:1974yy , their masses and interactions with fermions remain puzzling. If the LQs are vector bosons, their masses originate from spontaneous gauge symmetry breaking and are their mass is protected by gauge symmetry. However, if PS unification is realized at the multi-TeV scale, the LQ-fermion coupling structure needs to respect an approximate $U(2)^{n}$ flavor symmetry to evade the bounds from processes involving light flavors, such as $K_{L}\rightarrow\mu e$ Barbieri:2011ci ; Isidori:2012ts ; Barbieri:2012uh ; Blankenburg:2012nx ; Faroughy:2020ina ; Antusch:2023shi . Additional vector-like Fermions and/or extended gauge groups are typically needed to achieve TeV-scale partial unification while preserving $U(2)^{n}$ at low energies DiLuzio:2017vat ; Bordone:2017bld ; Greljo:2018tuh ; Calibbi:2017qbu ; Fuentes-Martin:2022xnb ; Davighi:2022bqf . On the other hand, scalar LQs can naturally preserve the chiral $U(2)^{n}$ symmetry, but their masses are unprotected and suffer from the same fine-tuning problem as the SM-Higgs boson PhysRevD.27.1601 . This conceptual puzzle becomes an explicit problem in a (partial) unification framework whose scale is much higher than a TeV. In our view, the LQ explanation for $B$ anomalies could become more convincing if LQs originate from a most minimal GUT. Their light masses and specific coupling structures could emerge as a consistency requirement for successful Yukawa unification, so that the benefits of successful unification and explanation of flavor anomalies outweigh the nuisance of additional unprotected scalar masses.
@@
The paper is structured as follows. In Section 2, we revisit the minimal ways to construct the SO(10) Yukawa sector. We explain why $126_{H}$ is preferred and discuss its shortcomings. Next, we discuss the LQ spectrum in $126_{H}$, specify the LQ-fermion interactions, and summarize the effective operators relevant for $b\rightarrow c,s$ transitions. In the following subsection, we present an overview on how these LQs can address the $B$ anomalies and examine what additional constrains are imposed by SO(10). In Section 3, we analyze how TeV-scale LQs can improve the RG evolution for the Yukawa couplings between light scalars and third-generation fermions. We show how the $b-\tau$ mass relationship is improved and how the LQ-fermion couplings exhibit fixed point behaviors. Since addressing $B$ anomalies requires large flavor mixing angles that are unphysical in SM, we then propose a possible solution to this newly arising problem. In Section 4, we summarize our main findings and outline the further work needed to strengthen our idea. To demonstrate the model-independence of our results, we show in Appendix A that $b-\tau$ unification can also be achieved with different light LQ spectra.
@@ 2.1 The landscape
Firstly, $126_{H}$ is tailor-made for tiny neutrino masses. $126_{H}$ contains a vacuum singlet under $G_{\text{SM}}$ and can give the right-handed neutrino $\nu_{R}$ a high-scale mass term. Upon integrating out $\nu_{R}$, the dim-5 Weinberg operator Weinberg:1979sa can generate a tiny mass for the active neutrino living in $\ell_{L}$ Minkowski:1977sc ; Mohapatra:1979ia . In addition, $126_{H}$ contains two Higgs doublets with opposite $SU(2)_{R}$ isospin. As a result, one of the Higgses only couples to $\overline{Q_{L}}u_{R},\overline{L_{L}}\nu_{R}$ and the other one only couples to $\overline{Q_{L}}b_{R},\overline{Q_{L}}e_{R}$. A hierarchical ratio $m_{t}/m_{b}$ is now possible, as long as the ratio $\tan\beta$ of the two Higgs doublets VEVs is large. Furthermore, $Y_{126}$ is symmetric and can be assigned third-generation specific, which gives the desired hierarchy structure among the charged fermion masses of the three generations. Indeed, Eq. (4) cannot account for quark mixing and is unlikely to yield the precise quantities for the first- and second-generation fermion masses. However, we do not think these shortcomings require extending the Yukawa sector immediately. If one neglects Yukawa couplings much smaller than unity, the light-flavor structure contains merely vanishing or unphysical observables. Large neutrino mixing is not problematic either, because the Yukawa couplings enter the left-handed neutrino mass matrix $M_{\nu_{L}}$ non-linearly. The relation $M_{\nu_{L}}\propto Y_{126}$ could break down even with higher-order corrections. Therefore, the $\mathcal{O}(1)$ predictions of $126_{H}$ are approximately correct. In the following, we refer to the theory whose Yukawa sector can be well-approximated by Eq. (4) as ‘minimal SO(10) GUT’.
@@ 2.2 Leptoquarks in $126_{H}$
Here,c denotes the standard charge conjugation operator, $Q_{R}=(u_{R},d_{R}),~L_{R}=(\ell_{R},\nu_{R})$ are $SU(2)_{R}$ doublets, and flavor indices are implicit. Eq. (8) shows $R$-type LQs couple to quarks and leptons with opposite chiralities, while the $S$-type LQs — $S_{3},S_{1}$ for left-handed fields and $\widehat{S}_{3},\widehat{S}_{1}$ for right-handed fields — are chirality-specific.
@@
$b\rightarrow c\tau\nu$ $(\overline{c}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma_{\mu}\nu_{L})$
@@
$(\overline{c}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma_{\mu}\nu_{L})$
@@
$b\rightarrow s\tau\tau$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{L}\gamma_{\mu}\tau_{L})$ - $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\tau}_{R}\gamma_{\mu}\tau_{R})$
$b\rightarrow s\nu\nu$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma_{\mu}\nu_{L})$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma_{\mu}\nu_{L})$ -
@@
$b\rightarrow s\tau\tau$ $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\tau}_{R}\gamma_{\mu}\tau_{R})$ $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\tau}_{L}\gamma_{\mu}\tau_{L})$
$b\rightarrow s\nu\nu$ - $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\nu}_{L}\gamma_{\mu}\nu_{L})$
@@
In the energy regime far below 1 TeV, LQs in Eq. (8) can be safely integrated out, and then the scalar-type effective operators with the form $(\overline{q}\ell)(\overline{\ell^{\prime}}q^{\prime})$ are generated at tree-level. We apply a Fierz transformation to convert these operators into the standard basis, as summarized in Table 1. These operators follow the general discussion shown in Ref. Dorsner:2016wpm and can also be directly inferred from the chirality structure of the LQ-fermion couplings. It is worth noting that the left-right symmetry is not manifest in the table, because $\nu_{R}$ and the top quark are not included. Moreover, the scalar- and tensor-type operators require sizable $R_{2}-(R_{2}^{\prime})^{c}$ and $S_{1}-(S_{1}^{\prime})^{c}$ (or $S_{1}-(S_{1}^{\prime\prime})^{c}$) mixing. Without such mixings, a single LQ in the generic basis cannot couple simultaneously to both $Q_{L}$ and $Q_{R}$, and thus leads to merely the vector-type operators.
@@ 2.3 Addressing the $B$ anomalies
TeV scale LQs are related to the long-standing anomalies observed in semi-leptonic B decays. The $S_{3}$ LQ contributes to all the $b\rightarrow c,s$ transitions with SM-like operators. If $S_{3}$ is the unique LQ responsible for the $R(D^{(*)})$ anomaly, $C_{L}^{\nu}$, the coefficient of $(\overline{s}_{L}\gamma^{\mu}b_{L})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ is fixed by $SU(2)_{L}$ invariance. The predicted value for $C_{L}^{\nu}$ is too large to be compatible with the current bounds on $b\rightarrow s\overline{\nu}\nu$ transitions. A simple way out is to combine $S_{3}$ with $S_{1}$, whose contributions to $C_{L}^{\nu}$ can partly cancel Crivellin:2017zlb ; Crivellin:2019dwb the one of the former LQ. Assuming a cancellation of about $60\%$ in $C_{L}^{\nu}$, and including the operators induced by $S_{1}$, $R(D^{(*)})$ can be consistently explained at $1\sigma$ level Crivellin:2025qsq . The constraints on $C_{L}^{\nu}$ can be relaxed since the recent Belle-II data with the inclusive tagging method indicates $\text{Br}(B\rightarrow K\overline{\nu}\nu)$ exceeds the SM prediction by a factor of 5.4. Yet $\text{Br}(B\rightarrow K^{*}\overline{\nu}\nu)$ still imposes a tight bound and necessitates the right-handed operator $(\overline{s}_{R}\gamma^{\mu}b_{R})(\overline{\nu}_{L}\gamma^{\mu}\nu_{L})$ generated by $\widetilde{R}_{2}$. If its coefficient $C_{R}^{\nu}$ takes a proper value, it can suppress the $B\rightarrow K^{*}$ amplitude while allowing a sizable $B\rightarrow K$ rate Bause:2023mfe ; He:2023bnk . $SU(2)_{L}$ invariance also implies $b\rightarrow s\overline{\tau}\tau$ transitions. The current limit on $\text{Br}(B\rightarrow K\overline{\tau}\tau)$ is weak due the experimental difficulty in identifying $\tau$ leptons. Interestingly, if closing the $\tau$ loop and attaching it to an off-shell photon, the penguin diagram can induce $b\rightarrow s\ell\ell$ ($\ell=e,\mu$) via lepton flavor universal operators Crivellin:2018yvo . The coefficient $C_{9}^{U}\sim-1$ can account for the $B\rightarrow K\ell\ell$ anomalies without violating the bound from $R(K^{(*)})$ Alguero:2023jeh . Although $B_{s}-\bar{B}_{s}$ mixing constraints disfavor the best-fit value for $C_{9}^{U}$, moderate cancellation by additional operators containing right-handed quarks can relax this tension Crivellin:2015era ; Capdevila:2023yhq . Such operators can be generated by $\widetilde{S}_{1}$ or $\widetilde{R}_{2}$.
The $R_{2}$ LQ can also explain the $R(D^{(*)})$ anomaly with the $(\overline{c}_{R}\sigma^{\nu\nu}b_{L})(\overline{\tau}_{R}\sigma_{\nu\nu}\nu_{L})$ and $(\overline{c_{R}}b_{L})(\overline{\tau}_{R}\nu_{L})$ operators Sakaki:2013bfa ; Becirevic:2018afm ; Cheung:2020sbq , and it does not lead to the $b\rightarrow s\overline{\nu}\nu$ transition. This solution requires an $\mathcal{O}(1)$ $R_{2}-(R_{2}^{\prime})^{c}$ mixing angle, which brings an inhomogeneous term to the RG equation for the $\tau$ lepton mass Fedele:2023rxb :
@@
The $S_{1}$ LQ is constrained to be as heavy as the GUT scale, because it always couples to a pair of quarks in the minimal GUT framework Patel:2022wya and induces proton decay. On the other hand, the diquark couplings of $S_{3},\overline{S}_{1},\widetilde{S}_{1}$ are absent at tree level. This absence is not accidental but a consequence of the $U(1)_{\text{PQ}}$ symmetry contained in Eq. (4).